Circular No. 8353

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004ch IN NGC 5612 R. Martin, Perth Observatory, reports the discovery of an apparent

supernova (red mag ∼ 16) found on CCD images taken on May 12.619 and June 1.431 UT with the 0.61-m Perth/Lowell Automated Telescope in the course of the Perth Automated Supernova Search. SN 2004ch is located at α = 14^{h}34^{m}06^{s}.27, δ = −78^{o}23′09′′.53 (equinox 2000.0; uncertainty ± 0′′.15 in each coordinate), which is 62′′.5 east and 8′′.0 north of the center of NGC 5612. Nothing was visible at this location on an image taken on May 12.619 (limiting mag 19).

SUPERNOVAE 2004bu, 2004bw, 2004bz, 2004cb, AND 2004cc T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Cen-

ter for Astrophysics, report that a spectrum (range 340–730 nm) of SN 2004cc (cf. IAUC 8350), obtained by K. Penev on June 10.46 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a supernova, most likely of type I, but the exact classification is uncertain. The spectrum appears to be highly reddened (a Na I D absorption at the host velocity has an equivalent width of 0.44 nm). There are peaks near 540 and 580 nm and absorptions near 560 and 630 nm.

R. J. Foley, D. S. Wong, M. Moore, and A. V. Filippenko, University of California, Berkeley, report that inspection of CCD spectra (range 330– 1000 nm), obtained on June 13 UT with the Shane 3-m telescope at Lick Observatory, shows that SN 2004bu (IAUC 8343) is similar to the peculiar type-Ic supernovae 1998bw and 2002ap (Foley et al. 2003, PASP 115, 1220), as previously suspected by Foley et al. (cf. IAUC 8345). SN 2004bw (IAUC 8345) is of type Ia, with a spectral-feature age of 7 ± 2 days after maximum brightness. SN 2004bz (IAUC 8350) and SN 2004cb in ESO 445-G20 (IAUC 8350) are also of type Ia, roughly 3 weeks past maximum brightness. The highly reddened SN 2004cc is of type Ic, probably about one week before maximum brightness.

POSSIBLE-SUPERNOVAE WEBPAGE A possible supernova found only recently on infrared images taken

of NGC 3256 in 1997 by S. Lipari et al. has been posted at the newly announced ‘CBAT Possible-Supernovae Page’ (cf. IAUC 8335), following referee recommendations, under the designation J9711-001 (which indicates its discovery in Nov. 1997). This highlights the fact that the page will be useful for various types of unconfirmed supernova candidates.

2004 June 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green